Without an appointment, the average waiting time in minutes at the doctor's office has the probability density function f(t)=1/38, where 0≤t≤38

Step 1 of 2:

What is the probability that you will wait at least 26 minutes? Enter your answer as an exact expression or rounded to 3 decimal places.

Step 2 of 2:

What is the average waiting time?
Answer #1
here for uniform distribution parameter a =0 and b=38

1)

probability = P(X>26)= 1-P(X<26)= 1-(26-0)/(38-0)= 0.316

2)

mean μ = (a+b)/2 = 19